OpenStudy (anonymous):

in 2008 the Florida marlins average attendance was 15,500. in 2009 the average attendance increased to 18,000. what was the percent of increase in the attendance to the marlins games ? Round your answer to the nearest tenth of a percent.

OpenStudy (lncognlto):

18,000 - 15,500 = 2500 Now 2500/15,500 = ?

OpenStudy (anonymous):

38,750000

OpenStudy (lncognlto):

O.o LOL No. \[2500 \div 15,500 = 0.1612903226\]

OpenStudy (lncognlto):

Now how do we convert 0.1612903226 into a percentage?

OpenStudy (anonymous):

it will be 16.1%

OpenStudy (lncognlto):

Yes, that is correct!

OpenStudy (anonymous):

thx i will be put a couple more questions

OpenStudy (lncognlto):

Okay xD Medal me maybe?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[percent increase =\frac{ 2500 }{15500 }*100\]

OpenStudy (lncognlto):

^^ Indeed.

OpenStudy (anonymous):

can u help sammyhhh

OpenStudy (anonymous):

she my cousin

OpenStudy (anonymous):

she on aipad and it will not let her write she has no figure it out yet

OpenStudy (anonymous):

i will try.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you have to wait . but you can post your question.

OpenStudy (anonymous):

ok